### v1

# Thresholded Lexicographic Ordered Multiobjective Reinforcement Learning

Abstract
Lexicographic multi-objective problems, which impose a lexicographic importance order over the objectives, arise in many real-life scenarios. Existing Reinforcement Learning work directly addressing lexicographic tasks has been scarce. The few proposed approaches were all noted to be heuristics without theoretical guarantees as the Bellman equation is not applicable to them. Additionally, the practical applicability of these prior approaches also suffers from various issues such as not being able to reach the goal state. While some of these issues have been known before, in this work we investigate further shortcomings, and propose fixes for improving practical performance in many cases. We also present a policy optimization approach using our Lexicographic Projection Optimization (LPO) algorithm that has the potential to address these theoretical and practical concerns. Finally, we demonstrate our proposed algorithms on benchmark problems.
[…]
2418
1 Introduction
The need for multi-objective reinforcement learning (MORL) arises in many real-life scenarios and the setting cannot be reduced to single-objective reinforcement learning tasks in general [22]. However, solving multiple objectives requires overcoming certain inherent difficulties. In order to compare candidate solutions, we need to incorporate given user preferences with respect to the different objectives. This can lead to Pareto optimal or non-inferior solutions forming a set of solutions where no solution is better than another in terms of all objectives. Various methods of specifying user preferences have been proposed and evaluated along three main fronts: (a)  expressive power, (b) ease of writing, and (c) the availability of methods for solving problems with such preferences. Typically there are tradeoffs when a particular front is emphasized. For example, writing preference specifications that result in a partial order of solutions instead of a total order makes the specification easier for the user but may not be enough to describe a unique preference. Three main motivating scenarios differing on when the user preference becomes available or used have been studied in the literature. (1) User preference is known beforehand and is incorporated into the problem a priori. (2) User preference is used a posteriori, i.e., firstly a set of representative Pareto optimal solutions is generated, and the user preference is specified over it. (3) An interactive setting where the user preference is specified gradually during the search and the search is guided accordingly.
Our present work is in the a priori setting. The most common specification method in this setting is linear scalarization which which requires the designer to assign weights to the objectives and take a weighted sum, thus making solutions comparable [2]. The main benefit of this technique is that it allows the use of many standard off the shelf algorithms as it preserves the additivity of the reward functions. However, expressing user preference with this technique requires significant domain knowledge and preliminary work in most scenarios [8]. While it can be the preferred method when the objectives can be expressed in comparable quantities, e.g. when all objectives have a monetary value, this is not the case most of the time. Usually, the objectives are expressed in incomparable quantities like money, time, and carbon emissions. Additionally, a composite utility combining the various objectives, and an approximation of that with linear scalarization limits us to a subset of the Pareto optimal set.
To address these drawbacks of linear scalarization, several other approaches have been proposed and studied. Nonlinear scalarization methods like Chebyshev [13] are more expressive and can capture all of the solutions in the Pareto optimal set, however, they do not address the user-friendliness requirement. In this paper, we will focus on an alternative specification method that overcomes both limitations of linear scalarization, named Thresholded Lexicographic Ordering (TLO) [3] [8]. In lexicographic ordering, the user determines an importance order for the objectives. The more important objectives dominate the lower order objectives. Given two candidate solutions, a lower ranked objective is used to rank the two solutions only if all higher order objectives have the same values. The thresholding part of the technique allows a more generalized definition for being the same w.r.t. an objective. The user provides a threshold for each objective except the last, and the objective values are clipped at the corresponding thresholds. This allows the user to specify values beyond which they are indifferent to the optimization of an objective. There is no threshold for the last objective as it is considered an unconstrained open-ended objective.
Despite the strengths of the TLO specification method, the need for a specialized algorithm to use it in reinforcement learning (RL) has prevented it from being a common technique. The Thresholded Lexicographic Q-Learning (TLQ) algorithm was proposed as such an algorithm and has been studied and used in several papers [8] [5]. While it has been noted that this algorithm does not enjoy the convergence guarantees of its origin algorithm (Q-Learning), we found that its practical use is limited to an extent that has not been discussed in the literature before. In this work, we investigate such issues of TLQ further. We also present a Policy Gradient algorithm as a general solution that has the potential to address many of the shortcomings of TLQ algorithms.
Our Contributions. Our main contributions in this work are as follows: (1) We demonstrate the shortcomings of existing TLQ variants on a new important problem class that was not known before. To the best of our knowledge, only Vamplew et al. [21] had previously shown a case where TLQ did not work. We formulate a taxonomy of the problem space in order to give insights into TLQ’s performance in different settings. Using our taxonomy, we demonstrate a new significant class of problems on which TLQ fails. We exhibit this new problematic class on a common control scenario where the primary objective is reaching a goal state and the other secondary objectives evaluate trajectories taken to the goal. (2) We propose a lexicographic projection algorithm which computes a lexicographically optimal direction that optimizes the current unsatisfied highest importance objective while preserving the values of more important objectives using projections onto hypercones of their gradients. Our algorithm allows adjusting how conservative the new direction is w.r.t. preserved objectives and can be combined with first-order optimization algorithms like Gradient Descent or Adam. We also validate this algorithm on a simple optimization problem from the literature. (3) We explain how this algorithm can be applied to policy-gradient algorithms to solve Lexicographic Markov Decision Processes (LMDPs) and experimentally demonstrate the performance of a REINFORCE adaptation on the cases that were problematic for TLQ.
2 Related Work
[3] was one of the first papers that investigate RL in multi-objective tasks with preference ordering. It introduces TLQ as an RL algorithm to solve such problems. Vamplew et al. [21] showed that TLQ significantly outperforms Linear Scalarization (LS) when the Pareto front is globally concave or when most of the solutions lie on the concave parts. However, LS performs better when the rewards are not restricted to terminal states, because TLQ cannot account for the already received rewards. Later, Roijers et al. [15] generalized this analysis by comparing more approaches using a unifying framework. To our knowledge, [21] is the only previous work that explicitly discussed shortcomings of TLQ. However, we found that TLQ has other significant issues that occur even outside of the problematic cases they analyze.
Wray et al. [25] introduced Lexicographic MDP (LMDP) and the Lexicographic Value Iteration (LVI) algorithm. LMDPs define the thresholds as slack variables which determine how worse than the optimal value is still sufficient. While Wray et al. [25] proved the convergence to desired policy if slacks are chosen appropriately, such slacks are generally too tight to allow user preferences. This is also observed in Pineda et al. [14] which claimed that while ignoring these slack bounds negates the theoretical guarantees, the resulting algorithm still can be useful in practice.
Li and Czarnecki [8] investigated the use of Deep TLQ for urban driving. It showed that the TLQ version proposed in [3] introduces additional bias which is especially problematic in function approximation settings like deep learning. Also, it depends on learning the true Q function, which can not be guaranteed. To overcome these drawbacks, it used slacks instead of the static thresholds and proposed a different update function. While Zhang et al. [28] also similarly utilize slacks, it defines them in terms of action probabilities instead of Q-values. Hayes et al. [5] used TLQ in a multi-objective multi-agent setting and proposed a dynamic thresholding heuristic to deal with the explosion of the number of thresholds to be set.
However, we discovered that these works on using a Q-learning variant with thresholded ordering perform very poorly in most cases due to non-Markovianity of the value function they try to learn. It is possible to bypass this issue by using policy gradient approaches as they do not require learning an optimal value function. In order to handle conflicting gradients, some modifications to the gradient descent algorithm are needed. Recent work on modified gradient descent algorithms came mostly from Multi Task Learning literature, which could be considered a multiobjective optimization problem [1] [17] [10] [12] [11]. While these papers use similar ideas with our work, their setting is different than ours as they do not have any explicit importance order; hence, not applicable to our setting. Uchibe and Doya [20] have the most similar setting to ours in gradient-based algorithms. It considers a set of constraints with an unconstrained objective. Then, the gradient of the unconstrained objective is projected onto the positive half-space of the active (violated) constraints and adds a correction step to improve the active constraints. When no valid projection is found, the most violated constraints are ignored until a valid projection exists. This is one of the main differences with our setting: As we have an explicit importance-order of the objectives, it is not acceptable to ignore a constraint without considering the importance order. Also, we project the gradients onto hypercones instead of hyperplanes, which is a hypercone with $\pi/2$ vertex angle. Thus, our algorithm allows varying degrees of conservative projections to prevent a decline in the constraints.
While there are many other recent works on Constrained Markov Decision Process (CMDPs) like [23, 6], their approaches are not applicable as an importance order over constraints is not allowed. Recently, using RL with lexicographic ordering began to attract attention from other communities as well. For example, Hahn et al. [4] use formal methods to construct single objective MDPs when all of the objectives are $\omega$-regular.
Skalse et al. [18] propose both value-based and policy-based approaches for LMDPs. Their value-based approach is based on slacks like [8] and they require using very small slack values. This protects their approach from the issues with having relaxations by limiting their setting to strict lexicographic order. For policy-based methods, they use Lagrangian relaxation and their setting is again a strict lexicographic ordering, i.e. it does not allow treating values above a threshold equal. As we are primarily interested in solving tasks with thresholded objectives, this paper does not address our question.
3 Background
Multiobjective Markov Decision Process (MOMDP). A MOMDP is a tuple $\langle S,A,P,\mathbf{R},\gamma\rangle$ where
  • $S$ is the finite set of states with initial state $s_{init}\in S$ and a set of terminal states $S_{F}$,
  • $A$ is a finite set of actions,
  • $P$: $S\times A\times S\to[0,1]$ is the state transition function given by $P(s,a,s^{\prime})=$ $\mathbb{P}(s^{\prime}|s,a)$, the probability of transitioning to state $s^{\prime}$ given current state $s$ and action $a$.
  • $\mathbf{R}=[R_{1},\dots,R_{K}]^{T}$ is a vector that specifies the reward of transitioning from state $s$ to $s^{\prime}$ upon taking action $a$ under $K$ different reward functions $R_{i}:S\times A\times S\to\mathbb{R}$ for $i\in\{1,\dots,K\}$.
  • $\gamma\in\mathbb{R}$ is a discount factor.
In such a MOMDP, a finite trajectory $\zeta\in(S\times A)^{*}\times S$ is a sequence $\zeta=s_{0}a_{0}s_{1}a_{1}\dots a_{T-1}s_{T}$ where $s_{i}\in S$, $a_{i}\in A$ and indices denote the time steps. The evolution of an MDP is governed by repeated agent-environment interactions, where in each step, an agent first picks actions in a state $s$ according to some probabilistic distribution, and for each of these actions $a$ the environment generates next states according to $\mathbb{P}(s^{\prime}|s,a)$. Each reward function $R_{i}$ corresponds to an objective $o_{i}$, the discounted rewards sum that the agent tries to maximize. Control over a MOMDP requires finding an optimal policy function $\pi^{*}:S\times A\to[0,1]$ which assigns probability $\mathbb{P}_{\pi^{*}}(a|s)$ to actions $a\in A$. In this paper, we use the episodic case of MDP where the agent-environment interaction consists of sequences that start in $s_{init}$ and terminates when a state in $S_{F}$ is visited. The length of the episode, T, is finite but not a fixed number. In MDP literature, this is known as "indefinite-horizon" MDP. The episodic case can be ensured by restricting ourselves to suitable policies which have a non-zero probability for all actions in all states.
We define the quality of a policy $\pi$ with respect to an objective $o_{i}\in\{1,\ldots,K\}$ by the value function $V^{\pi}_{i}:S\to\mathbb{R}$ given by $V^{\pi}_{i}(s)=\mathop{\mathbb{E}}_{\pi}[\sum_{t=0}^{T}\gamma^{t}R_{i}(s_{t},a_{t},s_{t+1})|s_{0}=s]$. Intuitively, $v^{\pi}_{i}(s)$ is the expected return from following policy $\pi$ starting from state $s$ w.r.t. objective $o_{i}$. The overall quality of a policy $\pi$ is given by the vector valued function $\boldsymbol{V^{\pi}}:S\to\mathbb{R}^{K}$ which is defined as $\boldsymbol{V^{\pi}}(s)=[V^{\pi}_{1}(s),\dots,V^{\pi}_{K}(s)]^{T}$. As $\boldsymbol{V}$ is vector-valued, without a preference for comparing $V^{\pi}_{i}$ values across different $i$, we only have a partial order over the range of $\boldsymbol{V}$, leading to Pareto front of equally good quality vectors. Further preference specification is needed to order the points on the Pareto front. A Lexicographic MDP (LMDP) is a class of MOMDP which provides such an ordering. It adds another component to MOMDP definition:
  • $\mathbf{\tau}=\langle\tau_{1},\ldots,\tau_{K-1}\rangle\in\mathbb{R}^{K-1}$ is a tuple of threshold values where $\mathbf{\tau}_{i}$ indicates the threshold value for objective $o_{i}$ beyond which there is no benefit for $o_{i}$. The last objective does not require a threshold; hence, there are only $K-1$ values. Then, $\boldsymbol{\tau}$ can be used to compare value vectors $\boldsymbol{u},\boldsymbol{v}\in\mathbb{R}^{K}$ by defining the thresholded lexicographic comparison $>^{\boldsymbol{\tau}}$ as $\mathbf{u}>^{\boldsymbol{\tau}}\mathbf{v}$ iff there exists $i\leq K$ such that:
    • $\forall j<i$ we have $\mathbf{u_{j}}\geq\min(\mathbf{v_{j}},\tau_{j})$; and
        if $i<K$ then $\min(\mathbf{u_{i}},\tau_{i})>\min(\mathbf{v_{i}},\tau_{i})$,
        otherwise if $i=K$ then $\mathbf{u_{i}}>\mathbf{v_{i}}$.
Intuitively, we compare $\boldsymbol{u}$ and $\boldsymbol{v}$ starting from the most important objective ($j=1$); the less important objectives are considered only if the order of higher priority objectives is respected.
Objectives $o_{i}$ for $i<K$ are said to be constrained objectives. A constrained objective $o_{i}$ is said to be satisfied when it is greater than or equal to its corresponding threshold $\tau_{i}$. The relation $\geq^{\boldsymbol{\tau}}$ is defined as $>^{\boldsymbol{\tau}}\textrm{ OR }=^{\boldsymbol{\tau}}$; where $\boldsymbol{u}=^{\boldsymbol{\tau}}\boldsymbol{v}$ if $\min(\mathbf{u_{i}},\tau_{i})=\min(\mathbf{v_{i}},\tau_{i})$ for each constrained objective and $\boldsymbol{u}_{K}=\boldsymbol{v}_{K}$.
Value-function Algorithms for Optimal Policies. An optimal policy $\pi^{*}$ is one that is better than or equal to any other policy, i.e., if $V^{\pi^{*}}(s)\geq^{\boldsymbol{\tau}}V^{\pi}(s)$ $\forall s\in S$ for all other policies $\pi\in\Pi$ [3]. There are two approaches to finding optimal policies in RL: Value-function algorithms and Policy Gradient algorithms. Value function based methods estimate the optimal action-value function $Q^{*}$ and construct $\pi^{*}$ using it. The action-value function under $\pi$, $Q^{\pi}:S\times A\to\mathbb{R}^{K}$, is defined as: $Q^{\pi}(s,a)\overset{\Delta}{=}\mathop{\mathbb{E}}_{\pi}[\sum_{t=0}^{T}\gamma^{t}\mathbf{R}(s_{t},a_{t},s_{t+1})|s_{0}=s,a_{0}=a]$. The optimal action-value function, $Q^{\star}$, is defined as: $Q^{\star}(s,a)=\max_{\pi\in\Pi}Q^{\pi}(s,a)$. Then, $\pi^{*}$ is obtained as: $\pi^{*}(s,a)=1$ if $a=\operatorname*{arg\,max}_{a^{\prime}\in A}Q^{\star}(s,a^{\prime})$, and 00 otherwise. In single objective MDPs, the Bellman Optimality Equation seen in Eq. 1 is used to learn $Q^{\star}$ as it gives an update rule that converges to $Q^{\star}$ when applied iteratively.
$Q^{\star}(s,a)=\mathop{\mathbb{E}}_{s^{\prime}\sim P}[(R(s,a,s^{\prime})+\gamma\max_{a^{\prime}\in A}Q^{\star}(s^{\prime},a^{\prime}))]$ (1)
Q-learning [24] is a very popular algorithm that takes this approach. TLQ tries to extend Q-learning for LMPDs; however, Bellman Optimality Equation does not hold in LMDPs. Hence, this approach lacks the theoretical guarantees enjoyed by Q-learning.
Policy Gradient Algorithms for $\pi^{*}$. Policy gradient algorithms in RL try to learn the policy directly instead of inferring it from the value functions. These methods estimate the gradient of the optimality measure w.r.t. policy and update the candidate policy using this potentially imperfect information. We denote the policy parameterized by a vector of variables, $\theta$ as $\pi_{\theta}$. The performance of the policy $\pi_{\theta}$, denoted $J(\theta)$, can be defined as the the expected return from following $\pi_{\theta}$ starting from $s_{init}$, i.e. $J(\theta)\overset{\Delta}{=}V^{\pi_{\theta}}(s_{init})$. Once the gradient of the optimality measure w.r.t. the parameters of the policy function is estimated, we can use first-order optimization techniques like Gradient Ascent to maximize the optimality measure. While all based on the similar theoretical results, a myriad of policy gradient algorithms have been proposed in the literature [19] [7] [16] [9].
4 TLQ: Value Function Based Approaches for TLO
Previous efforts to find solutions to the LMDPs have been focused on value-function methods. Apart from [25], which takes a dynamic programming approach, these have been variants of Thresholded Lexicographic Q-Learning (TLQ), an LMDP adaptation of Q-learning [3] [8]. While these methods have been used and investigated in numerous papers, the extent of their weaknesses has not been discussed explicitly.
In order to adapt Q-learning to work with LMDPs, one cannot simply use the update rule in Q-learning for each objective and learn the optimal value function of each objective completely independent of the others. Such an approach would result in the actions being suboptimal for some of the objectives. Based on this observation, two variants of TLQ [3][8] have been proposed, which differ in how they take the other objectives into account. We analyze these variants by dividing their approaches into two components: (1) value functions and update rules; and (2) acceptable policies for action selection. However, it should be noted that these components are inherently intertwined due to the nature of the problem — the value functions and acceptable policies are defined recursively where each of them uses the other’s last level of recursion. Due to these inherent circular referencing, the components will not be completely independent and some combinations of introduced techniques may not work.
Value Functions and Update Rules. The starting point of learning the action-value function for both variants is $Q^{\star}=\langle Q^{\star}_{1},\ldots,Q^{\star}_{K}\rangle$ where each $Q^{\star}_{i}:S\times A\to\mathbb{R}$ is defined as in Section 3 only with the change that the maximization is not done over the set of all policies $\Pi$ but over a subset of it $\Pi_{i-1}\subseteq\Pi$ which will be described below. Gábor et al. [3] propose learning $\hat{Q}^{\star}:S\times A\to\mathbb{R}^{K}$ where each component of $\hat{Q}^{\star}$ denoted by $\hat{Q}^{\star}_{i}$ is defined as: $\hat{Q}^{\star}_{i}(s,a)\overset{\Delta}{=}\min(\tau_{i},Q^{\star}_{i}(s,a))$ In other words, it is the rectified version of $Q^{\star}_{i}$. It is proposed to be learned by updating $\hat{Q}^{\star}_{i}(s,a)$ with the following value iteration which is adapted from Eq. 1
$\min\Big{(}\tau_{i},\sum_{s^{\prime}}P(s,a,s^{\prime})(R_{i}(s,a,s^{\prime})+\gamma\max_{\pi\in\Pi_{i-1}}\hat{Q}^{\star}_{i}\left(s^{\prime},\pi\left(s^{\prime}\right)\right)\Big{)}$ (2)
Notice that similar to the definition of $Q^{\star}_{i}$, the main change during the adaptation is limiting the domain of $\max$ operator to $\Pi_{i-1}$ from $\Pi$. On the other hand, Li and Czarnecki [8] propose that we estimate $Q^{\star}$ instead and use it when the actions are being picked. This $Q^{\star}$ uses the same update rule as Eq. 1 with only change being maximization over $\Pi_{i-1}$.
[…]
Function ActionSelection($s,Q,\epsilon|A$):
       $r\sim U(0,1)$
if $r<\epsilon$ then
             $a$ is picked randomly from $A$
       else
             $A_{0}\leftarrow A$
for o $=1,K$ do
                   if $|A_{o-1}|>1$ and $o<K$ then
[…]
                   else
                         $a\leftarrow\operatorname*{arg\,max}_{a^{\prime}\in A_{o-1}}Q_{i}(s,a^{\prime})$
break
             end for
       return $a$
Algorithm 1 ActionSelection
[…]
Absolute Thresholding: Gábor et al. [3] propose this approach where the actions with values higher than a real number are considered acceptable. Hence, $\Pi_{i}$ is the subset of $\Pi_{i-1}$ for which Q-values are higher than some $\tau$.
Absolute Slacking: This is the approach taken by Li and Czarnecki [8] where a slack from the optimal value in that state is determined and each action within that slack is considered acceptable.
4.1 Shortcomings of Prior TLQ Approaches
The shortcomings of TLQ depend on the type of task. We introduce a taxonomy to facilitate the discussion.
Constrained/Unconstrained Objective: Constrained objectives are bounded in quality by their threshold values above which all values are considered the same. Unconstrained objectives do not have a such threshold value. In an LMDP setting, all objectives but the last one are constrained objectives. Recall that a constrained objective is said to be satisfied when it exceeds its threshold.
Terminating Objective: An objective that either implicitly or explicitly pushes the agent to go to a terminal state of the MDP. More formally, this means discounted cumulative reward for an episode that does not terminate within the horizon is lower than an episode that terminates.
Reachability Objective: A reachability objective is represented by a non-zero reward in the terminal states of the MDP and zero rewards elsewhere. We will call an objective that has non-zero rewards in at least one non-terminal state a non-reachability objective.
This taxonomy also helps seeing the problematic cases and potential solutions. To summarize the empirical demonstrations about the applicability of TLQ for different parts of the problem space:
(I) TLQ successful case studies: Vamplew et al. [21] show an experiment where the constrained objective is reachability, unconstrained objective is terminating, and TLQ works. Li and Czarnecki [8] show this for a case study where the constrained objective is a non-terminating reachability constraint. Note that these are just demonstrations of empirical performance for case studies that fall into these categories and the papers do not make claims about general instances.
(II) TLQ does not work: Vamplew et al. [21] show that TLQ does not work when constrained objective is non-reachability. To the best of our knowledge, this is the only failure case that has been shown previously in the literature.
Proposition 4.1
In addition to previously identified failure case in Vamplew et al. [21], TLQ does not work when the constrained objective is a terminating reachability objective but the unconstrained one is non-terminating.
We next demonstrate Proposition 4.1’s claim on a maze problem.
MAZE Benchmark. We describe scenarios that are common in control tasks yet TLQ fails to solve. To illustrate different issues caused by TLQ, we need an adaptable multiobjective task. Also, limiting ourselves to finite state and action spaces where the tabular version of TLQ can be used simplifies the analysis. Our $\mathrm{MAZE}$ environment satisfies all of our requirements. Figure 1 shows an example.
|__|G_|__| 2 |HH|HH|__| 1 |__|S_|__| 0 0 1 2
Figure 1: A simple maze that demonstrates how TLQ fails to reach the goal state.
[…]
Problems with Reachability Constraint. A common scenario in control tasks is having a primary objective to reach a goal state and a secondary objective that evaluates the path taken to there. Formally, this is a scenario where the primary objective is a reachability objective. However, TLQ either needs to ignore the secondary objective or fails to satisfy the primary objective in such cases. All of the thresholding methods above fail to guarantee to reach the goal in this setting when used threshold/slack values are uniform throughout state space. The maze in Figure 1 can be used to observe this phenomenon. Assume that our primary objective is to reach $G$ and we encode this with a reward function that is 00 everywhere but $G$ where it is $R$. And our secondary objective is to avoid the bad tiles. A Pareto optimal policy in this maze could be indicated by the state trajectory $(1,0)\to(2,0)\to(2,1)\to(2,2)\to(1,2)$. However, this is unattainable by TLQ.
Since the reward is given only in the goal state, $\hat{Q}^{\star}_{1}$ can be equal to $\tau_{1}$ only for the state-action pairs that lead to the goal state. If no discounting was used, all actions would have the value $\tau_{1}$, hence the agent would not need to reach the goal state. We believe the reason why Vamplew et al. [21] have not observed this issue in their experiments with undiscounted ($\gamma=1$) Deep Sea Treasure (DST) is due to their objectives. The secondary objective of DST, minimizing the time steps before terminal state, is a terminating objective which pushes the agent to actually reach a goal state. Absolute Slacking also similarly leads to contradicting requirements when the trajectory described above is aimed but we leave this case to the reader due to space constraints.
5 Policy Gradient Approach for TLO
In this section, we introduce our policy gradient approach that utilizes consecutive gradient projections to solve LMDPs. Policy gradient methods treat the performance of a policy, $J(\theta)$, as a function of policy parameters that needs to be maximized and employ standard gradient ascent optimization algorithms. Following this modularity, we start with proposing a general optimization algorithm, the Lexicographic Projection Algorithm (LPA), for multiobjective optimization (MOO) problems with thresholded lexicographic objectives. Then, we will show how single objective Policy Gradient algorithms can be adapted to this optimization algorithm.
As gradients w.r.t. different objectives can be conflicting for MOO, various ways to combine them have been proposed. Uchibe and Doya [20] proposes projecting gradients of the less important objectives onto positive halfspaces of the more important gradients. Such a projection vector has the highest directional derivative w.r.t. the less important objective among directions with non-negative derivatives w.r.t. the important objectives. This is assumed to protect the current level for the important objective while trying to improve the less important. However, a non-negative derivative actually does not guarantee protection of the current level as infinitely small step sizes are not used in practice. For example, for a strictly concave function, the change in a zero-derivative direction will be always negative for any step size greater than 00. Therefore, we propose projecting onto hypercones which allows more control over how safe the projection is. A hypercone is the set of vectors that make at most a $\frac{\pi}{2}-\Delta$ angle with the axis vector; a positive halfspace is a special hypercone where $\Delta=0$. Increasing $\Delta$ brings the projection closer to the axis vector. A hypercone $C^{\Delta}_{a}$ with axis $a\in\mathbb{R}^{n}$ and angle $\frac{\pi}{2}-\Delta$ is defined as
$C^{\Delta}_{a}=\ \left\{\boldsymbol{x}\in\mathbb{R}^{n}\left\lvert\lVert\boldsymbol{x}\rVert=0\lor\frac{\boldsymbol{a}^{T}\boldsymbol{x}}{\lVert\boldsymbol{a}\rVert\lVert\boldsymbol{x}\rVert}\geq\cos{(\frac{\pi}{2}-\Delta)}\right.\right\}$ (3)
We can derive the equation for projection of $\boldsymbol{g}$ on $C^{\Delta}_{a}$ by firstly showing that $\boldsymbol{g}$, $\boldsymbol{a}$, and the projection $\boldsymbol{g^{p}}$ are planar using Karush-Kuhn-Tucker (KKT) conditions. Then, we can derive the formula by using two-dimensional geometry giving us $\boldsymbol{g^{p}}=$
$\!\frac{\cos{\Delta}}{\sin{\phi}}\sin{(\Delta+\phi)}\left(\!\boldsymbol{g}+\boldsymbol{a}\frac{\lVert\boldsymbol{g}\rVert}{\lVert\boldsymbol{a}\rVert}(\sin{\phi}\tan{\Delta}-\cos{\phi})\!\right)$ (4)
where $\phi$ is the angle between $\boldsymbol{a}$ and $\boldsymbol{g}$. Moving forward, we assume a function $projectCone(\boldsymbol{g},\boldsymbol{a},\Delta)$ which returns $\boldsymbol{g^{p}}$ according to this equation. Note that when $\boldsymbol{g}\in C^{\Delta}_{a}$, $projectCone(\boldsymbol{g},\boldsymbol{a},\Delta)\vcentcolon=\boldsymbol{g}$.
Lexicographic Projection Algorithm (LPA). A Thresholded Lexicographic MOO (TLMOO) problem with $K$ objectives and $n$ parameters can be formulated as maximizing a function $F:A\to\mathbb{R}^{K}$ where $A\subseteq\mathbb{R}^{n}$, and the ordering between two function values $F(\theta_{1}),F(\theta_{2})$ is according to $\geq^{\tau}$ as defined as in Section 3. Notice that when we have multiple objectives, the gradients will form a $K$-tuple, $G=(\nabla F_{1},\nabla F_{2},\cdots,\nabla F_{K})$, where $\nabla F_{i}$ is the gradient of $i^{th}$ component of $F$.
Since TLMOO problems impose a strict importance order on the objectives and it is not known how many objectives can be satisfied simultaneously beforehand, a natural approach is to optimize the objectives one-by-one until they reach the threshold values. However, once an objective is satisfied, optimizing the next objective could have a detrimental effect on the satisfied objective. We can use hypercone projection to avoid this. More formally, when optimizing $F_{i}$, we can project $\nabla F_{i}$ onto the intersection of $\{C^{\Delta}_{\nabla F_{j}}\}_{j<i}$ where $\Delta$ is a hyperparameter representing how risk-averse our projection is, and use the projection as the new direction. If such an intersection does not exist, it means that it is not possible to improve $F_{i}$ without taking a greater risk and we can terminate the algorithm.
We show in Proposition 5.1 that the LPA algorithm provides a strict increase in the value of the objective it currently optimizes while not decreasing the value of previously optimized objectives until it reaches a point where it cannot improve conservatively.
Proposition 5.1
Assume that when applied with a small enough step size on a TLMOO of $F_{i}$ concave and $L$-smooth for every $i\in\{1\dots K\}$, LPA yields the sequence in parameter space $\{\boldsymbol{x}^{(l)}\}_{l\in\mathbb{N}}$ with $\boldsymbol{x}^{(l)}\in\mathbb{R}^{n}\;\forall l\in\mathbb{N}$. Define the objective under optimization at step $l$ as $i^{*}_{l}=\min_{j}\{j<K:F_{j}(\boldsymbol{x}^{(l)})<\tau_{j}\}$ where $\min\{\}=K$. Then, $F_{j}(\boldsymbol{x}^{(l+1)})\geq F_{j}(\boldsymbol{x}^{(l)})$ for all $j\leq i^{*}_{l}$ and the relation is strict for $j=i^{*}_{l}$ when $\boldsymbol{x}^{(l)}$ is not the optimal point of $F_{i^{*}_{l}}$ or is a $\Delta$-Pareto-stationary point for the objective functions ${F_{1},F_{2},\dots,F_{i^{*}_{l}}}$.
The proof and the exact lower bounds on $F_{j}(\boldsymbol{x}^{(l+1)})-F_{j}(\boldsymbol{x}^{(l)})$ in the proposition are in the supplementary material.
Proposition 5.1 uses the concept of $\Delta$-Pareto-stationary points — a generalization of Pareto-stationary points that allows for parameterized conservativeness. A point $x$ is said to be $\Delta$-Pareto-stationary for the objective functions $\{F_{1},F_{2},\dots,F_{k}\}$ if there is not an ascent direction $\boldsymbol{u}$ such that $\max_{j\in{1\dots k}}\angle(\boldsymbol{u},\nabla F_{j})\leq\frac{\pi}{2}-\Delta$.
Remark 5.1
Under the $\mu$-strong convexity assumptions, a $\Delta$-Pareto-stationary point can be understood as a point where the norm of largest ascent step cannot be greater than $\frac{2\sin(\Delta)}{\mu}\max_{j}\lVert\nabla F_{j}(x)\rVert$.
Improving this approach with a heuristic that prevents overly conservative solutions leads to better performance in certain cases. Conservative updates usually lead to further increases on the already satisfied objectives instead of keeping them at the same level. This means most of the time, we have enough buffer between the current value of the satisfied objectives and their thresholds to sacrifice some of it for further gains in the currently optimized objective. Then, we can define a set of "active constraints" – a subset of all satisfied objectives – for which we will not accept any sacrifice and only consider these when projecting the gradient. The "active constraints" can be defined loosely, potentially allowing a hyperparameter that determines the minimum buffer needed to sacrifice from an objective.
Function FindDirection($M,F(\theta),\tau,\Delta,AC,b$):
       Initialize action-value function $Q$ with random weights
for o $=1,K$ do
             if $o=K$ or $F_{o}(\theta)<\tau_{o}$ then
                   Initialize direction $\boldsymbol{u}$ with initial state $M_{o}$
for j $=1,o-1$ do
                         if not (($AC$ and $F_{j}(\theta)>\tau_{j}+b$) or $\angle(\boldsymbol{u},M_{j})<\frac{\pi}{2}-\Delta$) then
                               $\boldsymbol{u}\leftarrow projectCone(\boldsymbol{u},M_{j},\Delta)$
                   end for
                  for j $=1,o$ do
                         if not (($j\neq K$ and $AC$ and $F_{j}(\theta)>\tau_{j}+b$) or $\angle(\boldsymbol{u},M_{j})<\frac{\pi}{2}-\Delta$) then
                               return None
                   end for
                  return $\boldsymbol{u}$
       end for
Algorithm 2 Lexicographic Constrained Ascent Direction
[…]
Experiment. As a benchmark for the Lexicographic Projection Algorithm, we used $F_{1}(x,y)=-4x^{2}+-y^{2}+xy$ and $F_{2}(x,y)=-(x-1)^{2}-(y-0.5)^{2}$ which are taken from [27]. We modified $F_{2}$ slightly for better visualization and multiplied both functions with $-1$ to convert this to a maximization problem. We set the threshold for $F_{1}$ to $-0.5$. The behavior of our cone algorithm without using active constraints heuristic on this problem with $\tau=(-0.5)$ and $\Delta=\frac{\pi}{90}$ can be seen in Figure 2.
Figure 2: The changes in the function values. Notice that $F_{2}$, in orange, is ignored until the threshold for $F_{1}$ is reached. Then, $F_{2}$ is optimized while respecting the passed threshold of $F_{1}$.
Using Lexicographic Projection Algorithm in RL. We show how LPA can be combined with policy gradient algorithms. We use REINFORCE[19] as the base policy gradient algorithm because its simplicity minimizes conceptual overhead.
[…]
       Initialize policy function $\pi(a|s,\theta)$ with random parameter $\theta$
for ep $=1,N_{e}$ do
             Generate an episode $S_{0},A_{0},R_{1},\ldots,S_{T-1},A_{T-1},R_{T}$ and save $\ln\pi(A_{t}|S_{t})$ at every step.
$M\leftarrow\emptyset$
$F\leftarrow 0$
for o $=1,K$ do
                   $G_{T+1}\leftarrow 0$
for t $=T,1$ do
                         $G_{t}\leftarrow R_{t}+\gamma G_{t+1}$
$F_{o}\leftarrow F_{o}+R_{t}$
                   end for
                  $L\leftarrow-\sum_{t=0,T-1}\ln\pi(A_{t}|S_{t})G_{t+1}$
Compute gradient of $L$ w.r.t. $\theta$ and append it to $M$
             end for
            $d=\textnormal{{FindDirection}}(M,F,\tau,\Delta,AC,b)$
Use $d$ as the gradient for the optimizer step to update $\theta$.
       end for
      return $\pi(a|s,\theta)$
Algorithm 3 Lexicographic REINFORCE
Note that our algorithm is compatible with most policy gradient algorithms. Uchibe and Doya [20] shows how a similar idea is applied to actor-critic family of policy gradient algorithms which reduces the variance in the gradient estimation by using a critic network. We believe that more stable policy gradient algorithms like actor-critic methods could further improve the performance of lexicographic projection approach as our algorithm might be sensitive to noise in gradient estimation.
6 Experiments
We evaluate the performance of the Lexicographic REINFORCE algorithm on two Maze problems and a many-objective benchmark from the literature. In both experiments, we use a two layer neural network for policy function.
|__|G_|__|__| 4 |__|hh|hh|hh| 3 |__|__|__|__| 2 |HH|HH|HH|__| 1 |S_|__|__|__| 0 0 1 2 3
Figure 3: The maze
Experiment 1: Non-reachability. The primary objective is non-reachability, i.e. it takes non-zero values in some non-terminal states. For this, we flip our objectives from the previous maze setting and define our primary objective as minimizing the cost incurred from the bad tiles. $HH$s give $-5$ reward and $hh$s give $-4$ reward. $+1$ reward is given in the terminal state to extend the primary objective to have rewards in both terminal and non-terminal states. The secondary objective is to minimize the time taken to the terminal state. We formalize this by defining our secondary reward function as 00 in terminal state and $-1$ everywhere else. We use the Maze in Figure 3 for this experiment. Note that this is the setting that Vamplew et al. [21] has found that TLQ fails. However, our experiments show that Lexicographic REINFORCE can solve this setting too.
Figure 4: Satisfaction rates for a single successful seed for the non-reachability maze experiment over 100 episodes.
We found that out of 10 seeds, 7 find policies that have 90% success over 100 episodes. The change in satisfaction rates of individual objectives for a successful seed can be seen in Figure 4. Notice that the primary objective initially starts very low and quickly increases while the secondary objective does the opposite. Once the primary objective is learned, the algorithm starts improving the secondary.
Experiment 2: Reachability (maze). As the second experiment, we consider the case where the primary objective is a reachability objective and the secondary objective is non-terminating, which was the setting that we found that TLQ fails to reach the goal state in Section 4.1. As the agent only cares about eventually reaching the goal, it can completely avoid going on a bad tile. We run Algorithm 3 for $N_{e}=4000$ episodes. Out of 10 seeds, 4 find policies that have 90% success over 100 episodes. Figure 5 shows how the satisfaction frequency for each objective changes throughout training.
Figure 5: Satisfaction rates for a single successful seed for the reachability maze experiment over 100 episodes.
These experiments illustrate the usefulness of projection based policy gradient algorithms for different tasks. We believe our results can be generalized to more complex tasks by combining our algorithm with more stable and sophisticated policy gradient algorithms.
Experiment 3: Reachability (FTN). Finally, we present our results on Fruit Tree Navigation (FTN) [26]. This task requires the agent to explore a full binary tree of depth $d=5$ with fruits on the leaf nodes. Each fruit has a randomly assigned vectorial reward $\mathbf{r}\in\mathbb{R}^{6}$ which encodes the amount of different nutrition components of the fruit. The agent needs to find a path from the root to the fruit that fits to the user preferences by choosing between left and right subtrees at every non-leaf node. As it has non-zero rewards only at the leafs, this is a reachability objective.
This domain perfectly highlights the benefits of thresholded lexicographic user preference compared to linear scalarization. The user can have a certain threshold that needs to be reached for each nutrition component and an importance order between these components that should be followed if it is not possible to satisfy all of the thresholds. Using linear scalarization in this case requires knowing the reward values of all of the fruits beforehand, deciding which fruit would fit the user preference, and finding a weight vector $\omega$ for which the desired fruit is better than others. Figure 6 compares our algorithm with TLQ on this domain. It also shows the need for the $\Delta$ parameter and highlights that using just hyperplanes as done in [20] fails this task. Similarly, TLQ agent also fails to find the desired leaf.
Figure 6: Probability of reaching the desired leaf, averaged over three different random seeds. $\Delta$ values represents conservativeness hyperparameter of the hypercone in degrees.
7 Conclusion
In this work, we considered the problem of solving LMDPs using model-free RL. While previous efforts on this problem have focused on value-function based approaches, applicability of these approaches over different problem settings and investigation of their shortcomings have been limited. Our first contribution was providing further insights into the inherent difficulties of developing value function based solutions to LMDPs. Towards this end, we demonstated a new significant class of failure scenarios for the existing methods.
Our second focus in this work was developing and presenting a policy-gradient based approach for LMDPs. Policy gradient algorithms have not been studied in MDPs with thresholded lexicographic objectives before even though they are more suitable for this setting as they bypass many inherent issues with value functions, such as non-convergence due to non-greedy policies w.r.t. value functions, and the need for different threshold values across the state space. For this, we developed a general thresholded lexicographic multi-objective optimization procedure based on gradient hypercone projections. Then, we showed how policy gradient algorithms can be adapted to work with LMDPs using our procedure, and demonstrated the performance of our REINFORCE adaptation on three case studies. While our results are promising for the REINFORCE adaptation, future research could be further empirical studies with more stable policy-gradient algorithm adaptations, and over more complex tasks.
{ack}
We thank Sarath Sreedharan, and the conference reviewers; their suggestions and feedback helped us improve the paper. This work was supported in part by the National Science Foundation by a CAREER award (grant number 2240126).